OpenStudy (anonymous):

Solve using the elimination method... 11x + 10y = 147 0x + 2y = 14

OpenStudy (araffi):

\[11x+10y=147\] \[0x+2y=14\] What you want to look for is a way to make the coefficients of either x or y the same (in absolute value). That way when you add or subtract the two equations, you will get a coefficient of 0 for that variable, which will allow you to solve for the other variable. The best way to do start is to find a variable where the coefficient in one equation is the multiple of another. In this case, 2-> one of the coefficients of y, multiplied by 5 = 10, the other coefficient of y. In order to keep the second equation true, you need to multiply both sides of the equation by 5. So the second equation becomes \[0x*5+2y*5=14*5\] OR \[0x+10y=70\] Subtracting the second equation from the first, you get 11x+0y=77 solving for x, x=77/11, so x=7, and plugging back into the original equation (or really just solving the second equation for y), you find that y is 7 also. [Hmmm, bizarre, is anyone else getting this answer?]